OpenStudy (anonymous):

equivalent forms of 2 1/4

OpenStudy (anonymous):

2*1/4

OpenStudy (anonymous):

\[2\tfrac{1}{4}=2+\frac{1}{4}=\frac{2\times 4+1}{4}=\frac{9}{4}\]